OpenStudy (asdfghj8):

WILL MEDAL FOR HELP! Two trains leave the station at the same time, one heading west and the other east. The westbound train travels 14 miles per hour slower than the eastbound train. If the two trains are 540 apart after 3 hours, what is the rate of the westbound train? Do not do any rounding.

OpenStudy (asdfghj8):

I know distance = rate * time but am unsure how to set up the equation here...

OpenStudy (otherworldly):

The eastbound train is traveling 180 miles per hr so

OpenStudy (otherworldly):

what do u think the rate is for the westbound train

OpenStudy (jiteshmeghwal9):

Let the speed of eastbound train be x the speed of westbound train = x+14 \[\frac{x}{3}+\frac{x+14}{3}=540\]

OpenStudy (jiteshmeghwal9):

x/3 is the distance covered by eastbound train in 3 hours (x+14)/3 is the distance covered by westbound train in 3 hours

OpenStudy (asdfghj8):

so \[\frac{ 2x+14 }{ 3 } = 540 \] ?

OpenStudy (asdfghj8):

x=803?????

OpenStudy (asdfghj8):

wow.... nope thats wrong.....not that i actually got any help...... its 83 if anyone looks it up